OpenStudy (anonymous):

Write as a single power of 2: (Rational indices) squareroot 2 ^3

OpenStudy (kinggeorge):

\[\sqrt{2^3} = 2^{3 \cdot {1 \over 2}} = 2^{3 \over 2}\]

OpenStudy (anonymous):

thankyou

OpenStudy (kinggeorge):

you're welcome